OpenStudy (anonymous):

the diameter of the moon is 3480km. (a) whats the surface area of the moon? (b)how many times larger is the surface area of the earth?

OpenStudy (anonymous):

surface area of a sphere is \[4\pi*r^2\] r=radius...

OpenStudy (anonymous):

assuming spherical shape to moon, its radius is 3480/2 km... substitute to find the area

OpenStudy (anonymous):

can u show me the work like how u got that

OpenStudy (anonymous):

got what?

OpenStudy (anonymous):

area?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

value of pi is approximately 3.14... just substitute surface area=4*(3.14)*(1740)^2 sq.km